# Generalized Bayesian Multidimensional Scaling and Model Comparison

### v2

@@ 3.2 Annealed SMC
where the forward kernel $\kappa^{+}(\boldsymbol{s}_{r-1,k},\boldsymbol{s}_{r,k})$ is a $\pi_{r}$-invariant Metropolis-Hastings kernel, and $\kappa^{-}(\boldsymbol{s}_{r,k},\boldsymbol{s}_{r-1,k})$ is the backward kernel (Del Moral et al., 2006). The selection of the backward kernel is crucial as it will affect the variance of the normalized weights. A convenient backward kernel that allows easy computation of the weight is
@@
$\tilde{w}_{r,k}=\left[\frac{l\left(\textbf{D}|\boldsymbol{s}_{r-1,k}\right)\pi(\boldsymbol{s}_{r-1,k})}{\tilde{\pi}_{0}(\boldsymbol{s}_{r-1,k})}\right]^{\tau_{r}-\tau_{r-1}}.$ (13)
@@
We sample the new particles $\boldsymbol{s}_{r,k}$ from $\pi_{r}$-invariant Metropolis-Hastings kernels. At each step $r$, we have the option to sample either all parameters or a subset of the parameters, depending on the desired trade-off between computational efficiency and effective exploration of the parameter space. In particular, for a high-dimensional parameter $\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{n}\right\}$, updating a subset of it can help mitigate the risk of low acceptance rates. This approach allows for more targeted exploration of the parameter space and potentially results in faster mixing. The annealed SMC algorithm can directly make use of the MCMC proposals in the particle propagation. The full conditional distributions for parameters $\lambda_{k}$, $\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{n}\right\}$, $\sigma^{2}$, and $\zeta_{i,j}$ are presented below. In each conditional posterior distribution, we use |$\cdots$| ⋯ to denote conditioning on the data and all other parameters and/or indicators. A detailed description of sampling methods is given in Supplementary.
@@
$\lambda_{k}|\cdots\sim IG(\alpha+n/2,\beta_{k}+\tau_{r}v_{k}/2),$ (14)
where $v_{k}/n$ is the sample variance of the $k$th coordinates of $\textbf{x}_{i}$’s.
The full conditional posterior distributions of $\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{n}\right\}$, $\sigma^{2}$ and $\psi$ do not admit closed forms, a random walk Metropolis-Hastings step is implemented with the Gaussian proposal densities.
@@
$\gamma_{r}\left(\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{n}\right\}|\cdots,\mathcal{M}_{TSN}\right)$ $\propto\exp\left\{-\tau_{r}\left(A+\frac{1}{2}\sum_{i=1}^{n}\textbf{x}_{i}^{\top}\Lambda^{-1}\textbf{x}_{i}\right)\right\},$
$\gamma_{r}\left(\sigma^{2}|\cdots,\mathcal{M}_{TSN}\right)$ $\propto\sigma^{-2(a+1)}\exp\left\{-\tau_{r}\left(A+\frac{b}{\sigma^{2}}\right)\right\},$
where $A=\frac{1}{2\sigma^{2}}\>SSR+\frac{m}{2}\log\left(\sigma^{2}\left(F_{\delta_{i,j},\sigma,\psi}(U)-F_{\delta_{i,j},\sigma,\psi}(0)\right)\right)-\sum\limits_{i>j}\log\left(\Phi\left(\psi\frac{d_{i,j}-\delta_{i,j}}{\sigma}\right)\right)$.
@@
$\gamma_{r}\left(\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{n}\right\}|\cdots,\mathcal{M}_{TT}\right)$ $\propto\exp\left\{-\tau_{r}\left(C+\frac{1}{2}\sum_{i=1}^{n}\textbf{x}_{i}^{\top}\Lambda^{-1}\textbf{x}_{i}\right)\right\},$
$\gamma_{r}\left(\sigma^{2}|\cdots,\mathcal{M}_{TT}\right)$ $\propto\sigma^{-m}\exp\left\{-\tau_{r}\left(C+\frac{b}{\sigma^{2}}\right)\right\},$
@@
$\zeta_{i,j}|\cdots,\mathcal{M}_{TT}\sim Gamma((\tau_{r}+\nu)/2,\tau_{r}(d_{i,j}-\delta_{i,j})^{2}/(2\sigma^{2})+\nu/4).$ (15)
@@
To alleviate the issue that all normalized weights converge to 0 except for one particle in sequential importance sampling, we prune particles of low weights when the population becomes too unbalanced. Popular resampling schemes include, but are not limited to, multinomial resampling, systematic resampling, stratified resampling, and residual resampling (Douc and Cappé, 2005). We will use multinomial resampling, as it is the most commonly used resampling method in the SMC literature.
Resampling at each iteration will increase the variance of the importance weights. Therefore, the resampling step is performed only when the degeneracy of the particles reaches some threshold $\epsilon$. At each iteration $r$, we monitor the degeneracy of particles using the effective sampling size (ESS) (Kong, 1992):
@@
The relative effective sample size (rESS) normalizes the ESS between zero and one. The rESS at iteration $r$ can be calculated by $\text{rESS}=\text{ESS}/K$.
@@
Input : (a) Initialization of $K$ particles: $\{\boldsymbol{s}_{0,k}\}_{k=1}^{K}$; (b) Priors and reference distributions over model parameters $\boldsymbol{s}=\{\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{n}\right\},\sigma^{2},\Lambda,\psi,\zeta\}$; (c) Likelihood function $l(\textbf{D}|\boldsymbol{s})$; (d) rCESS threshold $\phi$; (e) Resampling threshold $\epsilon$.
Output : (a) Particle population: $\{(\boldsymbol{s}_{R,k},W_{R,k})\}_{k=1}^{K}$; (b) Marginal likelihood estimates: $\widehat{M}_{R}$; (c) Total SMC iterations: R; (d) Sequence of annealing parameter: $\{\tau_{r}\}_{r=0}^{R}$.
@@
3Initialize SMC iteration index: $r\leftarrow 1$, initialize annealing parameter: $\tau_{0}\leftarrow 0$, initialize marginal likelihood estimate: $\widehat{M}_{0}\leftarrow 1$, load initial particles $\{\boldsymbol{s}_{0,k}\}_{k=1}^{K}$.
4for $k\in\{1,2,\ldots,K\}$ do
@@
9             Compute incremental importance weights: $\tilde{w}_{r,k}=\left[\frac{l\left(\textbf{D}|\boldsymbol{s}_{r-1,k}\right)\pi(\boldsymbol{s}_{r-1,k})}{\tilde{\pi}(\boldsymbol{s}_{r-1,k})}\right]^{\tau-\tau_{r-1}}.$
10      Determine the next annealing parameter $\tau_{r}$ using bisection method with: $f(\tau)=\text{rCESS}_{r}(W_{r-1,\cdot},w_{r,\cdot})=\phi$.
@@
12             Compute pre-resampling unnormalized weights: $w_{r,k}=w_{r-1,k}\times\tilde{w}_{r,k}$.
@@
14            Sample particles $\boldsymbol{s}_{r,k}$ from $\pi_{r}$-invariant Metropolis-Hastings kernels.
15
16      Update marginal likelihood estimates $\widehat{M}_{r}=\widehat{M}_{r-1}\times\sum^{K}_{k=1}W_{r-1,k}{\tilde{w}}_{r,k}$.
17      if $\tau_{r}=1$ then
@@
19            return $R$, $\{\tau_{r}\}_{r=0}^{R}$, $\{(\boldsymbol{s}_{R,k},W_{R,k})\}_{k=1}^{K}$, and $\widehat{M}_{R}$,
@@
21             if particle degeneracy is too severe, i.e. rESS $<\epsilon$ then
22                   Resample the particles, denoted $\{\boldsymbol{s}_{r,k}\}_{k=1}^{K}$; Reset particle weights: $w_{r,k}=1$, $W_{r,k}=1/K$.
23
24
@@
The annealed SMC algorithm produces a set of particles. After the extra resampling step in the end, the output of the annealed SMC algorithm contains a list of $K$ particles with equal weight. These particles can be used for the posterior approximation and for constructing the visualization in the lower-dimensional space. To find the Bayesian estimate of X, we take an approximate posterior mode of $\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{n}\right\}$ as described in Oh and Raftery (2001). Oh and Raftery (2001) observed that the term involving $SSR$ dominates the posterior density. Thus, the approximate posterior mode can be found by the values of $\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{n}\right\}$ that minimizes $SSR$ among all $K$ particles. The approximate posterior mode retrieves the relative positions of $\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{n}\right\}$, and they can be considered as the solution to the object configuration. Meaningful absolute positions of X may be obtained from some suitable transformation defined by the users if needed.
Some challenges arise with MCMC-based approximations in the context of model comparison via marginal likelihood estimators, as discussed in 2.3. MCMC-based algorithms often require additional computational costs to estimate the marginal likelihood separately. In contrast, model selection can be accomplished effortlessly using the Bayes factor in the proposed annealed SMC algorithm. When resampling is not performed at every step, the estimated marginal likelihood could be evaluated during the sampling process with the following formula:
$\widehat{M}_{R}=\prod_{r=1}^{R}\sum_{k=1}^{K}W_{r-1,k}\tilde{w}_{r,k}.$
Moreover, the estimates are unbiased when using a fixed annealing sequence and nearly unbiased with an adaptive annealing scheme. Previous studies have shown the advantages of SMC over MCMC for model comparison via marginal likelihood estimators (Zhou et al., 2016; Wang et al., 2020).
@@
We apply the adaptive annealing parameter scheme discussed in Wang et al. (2020). The main idea is to select an annealing parameter $\tau$ such that we achieve a controlled increase in particle degeneracy. The particle degeneracy between two successive intermediate distributions is measured by the relative conditional effective sample size (rCESS) (Zhou et al., 2016),
@@ 3.3 Adaptive mechanism
In the previous subsection, we presented the annealed SMC algorithm for a fixed dimension. In this section, we will describe an adaptive mechanism to enable the annealed SMC algorithm to handle increasing dimensions. The complete algorithm is presented in Algorithm 3.
Inferences are made sequentially for each batch of observations. When a new batch is available, the dissimilarity matrix $\textbf{D}=\{\textbf{D}^{(0)},\textbf{D}^{(1)}\}$ is calculated based on all the old observations and the incremental observations. The primary objective is to conduct inference for $\pi(\textbf{X}^{(0)},\textbf{X}^{(1)},\boldsymbol{\theta}|\textbf{D})$. In order to achieve this objective, we employ the strategy detailed in Section 3.1 for initializing the particles of $\boldsymbol{s}=\{\textbf{X}^{(0)},\textbf{X}^{(1)},\boldsymbol{\theta}\}$.
When previous estimations are unavailable, i.e., $\textbf{D}^{(0)}=\emptyset$, $\textbf{X}^{(0)}=\emptyset$, the reference distribution for $\textbf{X}^{(1)}=\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{n}\right\}$ is based on the results from fitting CMDS:
$\textbf{x}_{i}\sim\mathcal{N}\left(\textbf{x}_{i}^{\text{CMDS}},0.01\boldsymbol{I}\right),\text{independently for }i=1,\ldots,n,$
where $\textbf{x}_{i}^{\text{CMDS}}$ is the result from fitting CMDS on $\textbf{D}^{(1)}$. The variance of the reference distribution is selected in a manner that induces the particles to concentrate around the CMDS outputs. With a small value of the annealing parameter $\tau_{r}$, the intermediate target distribution is closer to the reference distribution, which is concentrated around the CMDS outputs for $\textbf{D}^{(1)}$.
When $\textbf{D}^{(0)}\neq\emptyset$ and $\textbf{X}^{(0)}\neq\emptyset$, we choose the reference distribution to be a particle approximation to the posterior distribution of $\textbf{X}^{(0)}$ given $\textbf{D}^{(0)}$. The reference distribution for $\textbf{X}^{(0)}=\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{n_{0}}\right\}$ is based on the results from all particles:
$\textbf{x}_{i}\sim\mathcal{N}\left(\hat{\textbf{x}}_{i},\hat{\boldsymbol{\Sigma}}\right),\text{independently for }i=1,\ldots,n_{0},$
where $\hat{\textbf{x}}_{i}$ is the particles’ posterior mode and $\hat{\boldsymbol{\Sigma}}$ is the estimated covariance matrix of all observations’ particles’ posterior modes from the previous computation. For the new incremental set $\textbf{X}^{(1)}$ in $\boldsymbol{s}$, we sample initial particles from the reference distribution, which is selected as its prior distribution for simplicity. This completes the specifications of the reference distributions in Algorithm 1. Initialization of particles is implemented from Line 3 to Line 3 of Algorithm 3.
@@
As an example, suppose we already obtain the posterior samples of $\textbf{X}^{(0)}$ from $n_{0}$ old observations by running the annealed SMC algorithm, and an extra of $n_{1}$ new observations become available. In that case, instead of running the annealed SMC algorithm from scratch using $n=n_{0}+n_{1}$ observations, we can utilize the information from the posterior samples of $\textbf{X}^{(0)}$ to initialize values for the old observations and use the prior distribution to initialize samples for the new observations $\textbf{X}^{(1)}$. One example that illustrates the case details with incremental dimensions is given in Section 5.1.
@@
In general, we can consider splitting data into $B$ batches; each batch has a size of $n_{b}-n_{b-1},$, for $b=1,\ldots,B$, with $n_{0}=0$. Figure 1 illustrates the setup for the adaptive mechanism when $b^{\text{th}}$ batch of data with size $n_{b}-n_{b-1}$ is observed after the posterior samples of $n_{b-1}$ old observations are obtained. Each batch is proceeded sequentially as outlined in Algorithm 3.
@@
Input : (a) Number of batches: $B$; (b) Data: $\textbf{Z}=\left\{\mathbf{z}_{1},\ldots,\mathbf{z}_{n}\right\}$; (c) Dissimilarity metric: $\mathcal{D}$.
Output : (a) Marginal likelihood estimates $\widehat{M}_{R}$; (b) Posterior approximation, $\hat{\pi}(\boldsymbol{s})=\sum^{K}_{k=1}W_{R,k}\times\delta_{\boldsymbol{s}_{R,k}}(\boldsymbol{s})$.
@@
8            Set $\tilde{\pi}_{1}\left(\textbf{x}_{i}\right)$ to $\mathcal{N}\left(\textbf{x}_{i}|\textbf{x}_{i}^{\text{CMDS}},0.01\boldsymbol{I}\right)$ for $i=1,\ldots,n_{1}$.
@@
10
11            Set $\tilde{\pi}_{0}\left(\textbf{x}_{i}\right)$ to $\mathcal{N}\left(\textbf{x}_{i}|\hat{\textbf{x}}_{i},\hat{\boldsymbol{\Sigma}}\right)$ for $i=1,\ldots,n_{b-1}$.
12            Set $\tilde{\pi}_{1}\left(\textbf{x}_{i}\right)$ to $\mathcal{N}\left(\textbf{x}_{i}|\textbf{0},\Lambda\right)$ for $i=n_{b-1}+1,\ldots,n_{b}$.
13
14      $\{\boldsymbol{s}_{0,k}\}_{k=1}^{K}\leftarrow$ Particle_Initialization $\left(\mathbf{D}^{(0)},\mathbf{D},\tilde{\pi}_{0}\left(\textbf{X}\right),\tilde{\pi}_{1}\left(\textbf{X}\right),\pi(\boldsymbol{\theta}),K\right)$
15      $\{(\boldsymbol{s}_{R,k},W_{R,k})\}_{k=1}^{K},\widehat{M}_{R}\leftarrow$ Annealed_SMC ($\left(\{\boldsymbol{s}_{0,k}\}_{k=1}^{K},\tilde{\pi}_{0}\left(\textbf{X}\right),\tilde{\pi}_{1}\left(\textbf{X}\right),\pi(\boldsymbol{\theta}),l(\textbf{D}|\boldsymbol{s}),\phi,\epsilon\right)$
16      Posterior approximation: $\hat{\pi}(\boldsymbol{s}^{(b)})=\sum^{K}_{k=1}W_{R,k}\times\delta_{\boldsymbol{s}_{R,k}}(\boldsymbol{s})$.
17      Compute the weighted mean $\hat{\textbf{x}}_{i}$ and covariance $\hat{\boldsymbol{\Sigma}}$, $i=1,\ldots,n_{b}$ from $\{(\boldsymbol{s}_{R,k},W_{R,k})\}_{k=1}^{K}$.

### v3

@@ 3.2 Annealed SMC
where the forward kernel $\kappa^{+}(\boldsymbol{s}_{r-1,k},\boldsymbol{s}_{r,k})$ is a $\pi_{r}$-invariant Metropolis-Hastings kernel, and $\kappa^{-}(\boldsymbol{s}_{r,k},\boldsymbol{s}_{r-1,k})$ is the backward kernel [del2006sequential]. The selection of the backward kernel is crucial as it will affect the variance of the normalized weights. A convenient backward kernel that allows easy computation of the weight is
@@
$\tilde{w}_{r,k}=\left[\frac{L\left(\boldsymbol{s}_{r-1,k};\textbf{D}\right)\pi(\boldsymbol{s}_{r-1,k})}{\tilde{\pi}_{0}(\boldsymbol{s}_{r-1,k})}\right]^{\tau_{r}-\tau_{r-1}}.$ (13)
@@
We sample the new particles $\boldsymbol{s}_{r,k}$ from $\pi_{r}$-invariant Metropolis-Hastings kernels. At each step $r$, we have the option to sample either all parameters or a subset of the parameters, depending on the desired trade-off between computational efficiency and effective exploration of the parameter space. In particular, for a high-dimensional parameter $\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{n}\right\}$, updating a subset of it can help mitigate the risk of low acceptance rates. This approach allows for more targeted exploration of the parameter space and potentially results in faster mixing. The full conditional distributions for latent variables $\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{n}\right\}$, and parameters $\lambda_{k}$, $\sigma^{2}$, and $\zeta_{i,j}$ are presented below. In each full conditional distribution, we use $|\cdot$ to denote conditioning on the data and all other parameters. A detailed description of sampling methods is given in Supplementary.
@@
$\lambda_{k}|\cdot\sim IG(\alpha+n/2,\beta_{k}+\tau_{r}v_{k}/2),$ (14)
where $v_{k}$ is the quantity for which $v_{k}/n$ is the sample variance of the $k$th coordinates of $\textbf{x}_{i}$’s.
The full conditional distributions of $\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{n}\right\}$, $\sigma^{2}$ and $\psi$ do not admit closed forms, a random walk Metropolis-Hastings step is implemented with the Gaussian proposal densities.
@@
$\gamma_{r}\left(\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{n}\right\}|\cdot\right)$ $\propto\exp\left\{-\tau_{r}\left(A+\frac{1}{2}\sum_{i=1}^{n}\textbf{x}_{i}^{\top}\Lambda^{-1}\textbf{x}_{i}\right)\right\},$ (15)
$\gamma_{r}\left(\sigma^{2}|\cdot\right)$ $\propto\sigma^{-2(a+1)}\exp\left\{-\tau_{r}\left(A+\frac{b}{\sigma^{2}}\right)\right\},$ (16)
where
$A=\frac{1}{2\sigma^{2}}\>SSR+\frac{m}{2}\log\left(\sigma^{2}\left(F_{\delta_{i,j},\sigma,\psi}(U)-F_{\delta_{i,j},\sigma,\psi}(0)\right)\right)-\sum\limits_{i>j}\log\left(\Phi\left(\psi\frac{d_{i,j}-\delta_{i,j}}{\sigma}\right)\right)$.
@@
$\gamma_{r}\left(\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{n}\right\}|\cdot\right)$ $\propto\exp\left\{-\tau_{r}\left(C+\frac{1}{2}\sum_{i=1}^{n}\textbf{x}_{i}^{\top}\Lambda^{-1}\textbf{x}_{i}\right)\right\},$ (17)
$\gamma_{r}\left(\sigma^{2}|\cdot\right)$ $\propto\sigma^{-m}\exp\left\{-\tau_{r}\left(C+\frac{b}{\sigma^{2}}\right)\right\},$ (18)
@@
$\zeta_{i,j}|\cdot\sim Gamma((\tau_{r}+\nu)/2,\tau_{r}(d_{i,j}-\delta_{i,j})^{2}/(2\sigma^{2})+\nu/4).$ (19)
@@
To alleviate the issue of particle degeneracy, where all normalized weights approach to 0 except for one, we resample the particles to replicate those with high weights and discard those with low weights. Popular resampling schemes include, but are not limited to, multinomial, systematic, stratified, and residual resampling [douc2005comparison]. We employ multinomial resampling. Although resampling alleviates particle degeneracy, performing it at every iteration introduces excess Monte Carlo variance into the estimator. Therefore, we perform resampling only when the degeneracy of the particles exceeds a specific threshold $\epsilon$. We monitor this using the effective sampling size (ESS) [kong1992note]:
@@
The relative effective sample size (rESS), defined as $\text{rESS}=\text{ESS}/K$, normalizes this metric to the interval [0,1].
@@
Input : (a) Initialization of $K$ particles: $\{\boldsymbol{s}_{0,k}\}_{k=1}^{K}$; (b) Priors $\pi(\boldsymbol{s})$ and reference distributions $\tilde{\pi}(\boldsymbol{s})$ for $\boldsymbol{s}=\{\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{n}\right\},\sigma^{2},\Lambda,\psi,\zeta\}$; (c) Likelihood function $L(\boldsymbol{s};\textbf{D})$; (d) rCESS threshold $\phi$; (e) Resampling threshold $\epsilon$.
Output : (a) Particle population: $\{(\boldsymbol{s}_{R,k},W_{R,k})\}_{k=1}^{K}$; (b) Marginal likelihood estimates: $\widehat{Z}_{R}$; (c) Total SMC iterations: $R$; (d) Sequence of annealing parameter: $\{\tau_{r}\}_{r=0}^{R}$.
@@
31ex
4Initialize SMC iteration index: $r\leftarrow 0$, initialize annealing parameter: $\tau_{0}\leftarrow 0$, initialize marginal likelihood estimate: $\widehat{Z}_{0}\leftarrow 1$, load initial particles $\{\boldsymbol{s}_{0,k}\}_{k=1}^{K}$.
5for $k\in\{1,2,\ldots,K\}$ do
@@
10     Express the incremental importance weights as a function of $\tau_{r}$: $\tilde{w}_{r,k}=\left[\frac{L\left(\boldsymbol{s}_{r-1,k};\textbf{D}\right)\pi(\boldsymbol{s}_{r-1,k})}{\tilde{\pi}(\boldsymbol{s}_{r-1,k})}\right]^{\tau_{r}-\tau_{r-1}}.$
11 Determine the next annealing parameter $\tau_{r}$ using bisection method with: $f(\tau)=\text{rCESS}_{r}(W_{r-1,\cdot},\tilde{w}_{r,\cdot})=\phi$.
@@
13     Compute unnormalized weights: $w_{r,k}=w_{r-1,k}\,\tilde{w}_{r,k}$.
@@
15    Sample particles $\boldsymbol{s}_{r,k}$ from $\pi_{r}$-invariant Metropolis-Hastings kernels using (14) -(19).
16
17 Update marginal likelihood estimates $\widehat{Z}_{r}=\widehat{Z}_{r-1}\,\sum^{K}_{k=1}W_{r-1,k}{\tilde{w}}_{r,k}$.
18 if $\tau_{r}\geq 1$ then
19     Set $\tau_{r}\leftarrow 1$. Repeat Steps 5–6 and 8–12.
@@
21    Return $R$, $\{\tau_{r}\}_{r=0}^{R}$, $\{(\boldsymbol{s}_{R,k},W_{R,k})\}_{k=1}^{K}$, and $\widehat{Z}_{R}$.
@@
23     if particle degeneracy is too severe, i.e. rESS $\epsilon$ then
24        Resample the particles, denoted $\{\boldsymbol{s}_{r,k}\}_{k=1}^{K}$;
25       1exReset particle weights: $w_{r,k}=1$, $W_{r,k}=1/K$.
26
27
@@ 24        Resample the particles, denoted $\{\boldsymbol{s}_{r,k}\}_{k=1}^{K}$;
The annealed SMC algorithm produces a set of particles. After the extra resampling step in the end, the output of the annealed SMC algorithm contains a list of $K$ particles with equal weight. These particles can be used for the posterior approximation and for constructing the visualization in the lower-dimensional space. To find the Bayesian estimate of X, we take an approximate posterior mode of $\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{n}\right\}$ as described in oh2001bayesian. oh2001bayesian observed that the term involving $SSR$ dominates the posterior density. Thus, the approximate posterior mode can be found by the values of $\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{n}\right\}$ that minimizes $SSR$ among all $K$ particles. The approximate posterior mode retrieves the relative positions of $\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{n}\right\}$, and they can be considered as the solution to the object configuration. Meaningful absolute positions of X may be obtained from some suitable transformation defined by the users if needed.
Some challenges arise with MCMC-based approximations in the context of model comparison via marginal likelihood estimators, as discussed in 2.3. MCMC-based algorithms often require additional computational costs or complex post-processing to estimate the marginal likelihood separately. In contrast, model selection can be accomplished directly using the Bayes factor in the proposed annealed SMC algorithm, where the marginal likelihood is a natural byproduct of the sampling process. The estimated marginal likelihood is evaluated using the following formula:
$\widehat{Z}_{R}=\prod_{r=1}^{R}\sum_{k=1}^{K}W_{r-1,k}\tilde{w}_{r,k}.$
Moreover, these estimates are unbiased when using a fixed annealing sequence and nearly unbiased with an adaptive annealing scheme. Previous studies have shown the advantages of SMC over MCMC for model comparison [zhou2016toward, wang2020annealed].
@@
We apply the adaptive annealing parameter scheme. The main idea is to select an annealing parameter $\tau$ such that we achieve a controlled increase in particle degeneracy. The particle degeneracy between two successive intermediate distributions is measured by the relative conditional effective sample size (rCESS) [zhou2016toward],
@@ 3.3 Adaptive Bayesian inference
We have presented the annealed SMC algorithm for a fixed dataset. In this section, we describe a sequential mechanism to enable the algorithm to handle incremental batches of data. The complete algorithm is presented in Algorithm 3.
@@
Suppose we have already obtained the posterior samples of $\textbf{X}^{(0)}$ from $m_{0}$ old data points by running the annealed SMC algorithm, and an additional $m_{1}$ new data points become available. In this case, instead of running the annealed SMC algorithm from scratch using the combined dataset of size $n=m_{0}+m_{1}$, we can utilize the information from the existing posterior samples to initialize the particles for $\textbf{X}^{(0)}$ and use the prior distribution to initialize samples for the new latent variables $\textbf{X}^{(1)}$.
More specifically, inferences are updated sequentially as new batches of data arrive. When a new batch becomes available, the dissimilarity matrix $\textbf{D}=\textbf{D}^{(0)}\cup\textbf{D}^{(1)}$ is calculated based on all available data including both old data of size $m_{0}$ and new data of size $m_{1}$. The objective is to conduct inference for the joint posterior $\pi(\textbf{X}^{(0)},\textbf{X}^{(1)},\boldsymbol{\theta}|\textbf{D})$. To achieve this, we employ the strategy detailed in Section 3.1 to initialize the particles of $\boldsymbol{s}=\{\textbf{X}^{(0)},\textbf{X}^{(1)},\boldsymbol{\theta}\}$.
When no previous data are available, we have $\textbf{D}^{(0)}=\emptyset$, $\textbf{X}^{(0)}=\emptyset$, $m_{0}=0$, and $n=m_{1}$. The reference distribution for the initial latent variables $\textbf{X}^{(1)}=\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{n}\right\}$ is constructed using the solution from CMDS:
$\textbf{x}_{i}\sim\mathcal{N}\left(\textbf{x}_{i}^{\text{CMDS}},\rho\boldsymbol{I}\right),\text{ for }i=1,\ldots,n,$
where $\textbf{x}_{i}^{\text{CMDS}}$ is the result from fitting CMDS on $\textbf{D}^{(1)}$. The variance $\rho$ is selected to concentrate particles around the CMDS solution. With a small value of the annealing parameter $\tau_{r}$, the intermediate target distributions remain close to this data-driven initialization.
When prior information exists ($\textbf{D}^{(0)}\neq\emptyset$ and $\textbf{X}^{(0)}\neq\emptyset$), we construct the reference distribution for the existing latent variables $\textbf{X}^{(0)}$ using a Gaussian approximation derived from the previous posterior samples. Specifically:
$\textbf{x}_{i}\sim\mathcal{N}\left(\hat{\textbf{x}}_{i},\hat{\boldsymbol{\Sigma}}\right),\text{ for }i=1,\ldots,m_{0},$
where $\hat{\textbf{x}}_{i}$ is the posterior mode of $\textbf{x}_{i}$ and $\hat{\boldsymbol{\Sigma}}$ represents the estimated covariance matrix computed from the previous particle set. For the new incremental set $\textbf{X}^{(1)}$ in $\boldsymbol{s}$, we sample initial particles from the reference distribution, which is chosen to be its prior distribution for simplicity. This completes the specifications of the reference distributions in Algorithm 1. Initialization of particles is implemented from Line 3 to Line 3 of Algorithm 3.
In general, we can consider splitting data into $B$ batches, where each batch $b$ (for $b=1,\ldots,B$) has a size $n_{b}-n_{b-1},$ with $n_{0}=0$. Figure 1 illustrates the setup for this adaptive mechanism where the $b^{\text{th}}$ batch is observed after the posterior samples for the accumulated data of size $n_{b-1}$ have been obtained. Each batch is processed sequentially as outlined in Algorithm 3. Section 5.3 provides an example to illustrate the details with sequential data updates.
@@
Input :  (a) Data $\textbf{Z}=\left\{\mathbf{z}_{1},\ldots,\mathbf{z}_{n}\right\}$ and cumulative batch sizes $n_{1},\ldots,n_{B}$; (b) Dissimilarity metric $\mathcal{D}$; (c) Tuning parameters $\phi,\epsilon$.
Output : (a) Marginal likelihood estimates $\widehat{Z}_{R}$; (b) Posterior approximation, $\hat{\pi}(\boldsymbol{s})=\sum^{K}_{k=1}W_{R,k}\,\delta_{\boldsymbol{s}_{R,k}}(\boldsymbol{s})$.
@@
7    Set $\tilde{\pi}_{1}\left(\textbf{x}_{i}\right)$ to $\mathcal{N}\left(\textbf{x}_{i}|\textbf{x}_{i}^{\text{CMDS}},\rho\boldsymbol{I}\right)$ for $i=1,\ldots,n_{1}$.
@@
9
[…]
11    Set $\tilde{\pi}_{1}\left(\textbf{x}_{i}\right)$ to the prior distribution $\mathcal{N}\left(\textbf{x}_{i}|\textbf{0},\Lambda\right)$ for $i=n_{b-1}+1,\ldots,n_{b}$.
12
13 $\{\boldsymbol{s}_{0,k}\}_{k=1}^{K}\leftarrow$ Particle_Initialization $\left(\mathbf{D}^{(0)},\mathbf{D},\tilde{\pi}_{0},\tilde{\pi}_{1},\pi(\boldsymbol{\theta}),K\right)$
14 $\{(\boldsymbol{s}_{R,k},W_{R,k})\}_{k=1}^{K},\widehat{Z}_{R}\leftarrow$ Annealed_SMC $\left(\{\boldsymbol{s}_{0,k}\}_{k=1}^{K},\tilde{\pi}_{0},\tilde{\pi}_{1},\pi(\boldsymbol{\theta}),L(\boldsymbol{s};\textbf{D}),\phi,\epsilon\right)$
15 Posterior approximation: $\hat{\pi}(\boldsymbol{s}^{(b)})=\sum^{K}_{k=1}W_{R,k}\,\delta_{\boldsymbol{s}_{R,k}}(\boldsymbol{s})$.
16 Compute weighted mean $\{\hat{\textbf{x}}_{i}\}_{i=1}^{n_{b}}$ and covariance $\hat{\boldsymbol{\Sigma}}$ from $\{(\boldsymbol{s}_{R,k},W_{R,k})\}_{k=1}^{K}$.